OpenStudy (anonymous):

HELP!!!!! ASAP!!!!!!! 2X + 1 = 5X - 13

OpenStudy (anonymous):

2x+1 = 5x-13 -5x -5x 2x-5x+1=-13 -1 -1

OpenStudy (anonymous):

2x-5x=-13-1

OpenStudy (anonymous):

2x-5x=-14

OpenStudy (anonymous):

-3x=-14

OpenStudy (anonymous):

x=4.67

OpenStudy (anonymous):

did you get it?